Find the vertex of the parabola.y = -4x2 - 16x - 11
Answer by stanbon(75887)

Find the vertex of the parabola.
y = -4x^2 - 16x - 11
---
Vertex occurs at x = -b/2a = 16/(-8) = -2
f(-2)= -4(-2)^2 - 16(-2) - 11
f(-2) = -16 + 32 - 11
f(-2) = 5
---
Vertex:: (-2,5)
--------------
Cheers,
Stan H.